OpenStudy (anonymous):

i really need help with algebra 2

OpenStudy (anonymous):

1. Find all the real fourth roots of . (1 point) and 4/7 and 16/49 -4/7 and 4/7 4/7,-4/7,16/49,,-16/49 4/7

OpenStudy (anonymous):

I got -4/7 amd 4/7 which is B

OpenStudy (anonymous):

2. What is a simpler form of the radical expression? (1 point)36|g3| 36g4 6|g3| 6g4

OpenStudy (anonymous):

lemme guess, the absolute value bars kinda threw you off?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well just ignore them, pretend the choices are: A.) 36g^3 B.) 36g^4 C.) 6g^3 D.) 6g^4

OpenStudy (anonymous):

then find the square root of 36 which is?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

good, then find x+x=6 for the exponent of the g

OpenStudy (anonymous):

is it c

OpenStudy (anonymous):

yepp thats right

OpenStudy (anonymous):

The formula for the volume of a sphere is v=3/4pie r ^3. Find the radius, to the nearest hundredth, of a sphere with a volume of 15 in.3 (1 point)

OpenStudy (anonymous):

Okay I had problems with this one too but I got it. Set it up like this: 15=4/3(3.14)r^3 ok?

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

then to divide by a fraction, you just multiply by its reciprocal like this: 15(3/4)=3.14r^3

OpenStudy (anonymous):

i multiply 15(3/4) together

OpenStudy (anonymous):

yeah then you'll have 45/4 which is 11.25. so then you have 11.25=3.14r^3

OpenStudy (anonymous):

do i now divide 11.25 with3.14

OpenStudy (anonymous):

yepp

OpenStudy (anonymous):

3.11r3

OpenStudy (anonymous):

no,11.25/3.14 should be 3.58 then 3.58=r^3 then find the cube root of 3.58 to get r.

OpenStudy (anonymous):

i got 3

OpenStudy (anonymous):

for the cube root of 3.58?

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

try it again

OpenStudy (anonymous):

i dont remember how to get the cube root

OpenStudy (anonymous):

\[\sqrt[3]{3.58}\]

OpenStudy (anonymous):

find x*x*x-3.58, it's a decimal. you'll get r=1.52977541 then round to get r=1.53 in.

OpenStudy (anonymous):

10. What is the simplest form of the quotient? (1 point)

OpenStudy (anonymous):

\[\sqrt[3]{\frac{ 162 }{ 2 }}\]

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

7. Kepler’s Third Law of Orbital Motion states that you can approximate the period P (in Earth years) it takes a planet to complete one orbit of the sun using the function , where d is the distance (in astronomical units, AU) from the planet to the sun. How many Earth years would it take for a planet that is 6.76 AU from the sun? (1 point)15.23 17.58 154.46 3.58

OpenStudy (anonymous):

no you have to simplify more, what is 162/2

OpenStudy (anonymous):

81

OpenStudy (anonymous):

good, now what is the square root of 81?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

okay so it goes 81 / \ 9 9 /\ /\ 3 3 3 3 now since your finding the cube root, you have to find 3 of the same number. are there any?

OpenStudy (anonymous):

yes three

OpenStudy (anonymous):

okay so move 3 to the outside of the radical like this\[3\sqrt[3]{x}\]

OpenStudy (anonymous):

but there's 4 three's so there is one left under the radical after you move the others out in front of the radical

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

yeah, so whats the final answer look like ?

OpenStudy (anonymous):

3 square root of x/3

OpenStudy (anonymous):

close, but it would be a cube root instead of a square root. so its \[3\sqrt[3]{3}\]

OpenStudy (anonymous):

7. Kepler’s Third Law of Orbital Motion states that you can approximate the period P (in Earth years) it takes a planet to complete one orbit of the sun using the function , where d is the distance (in astronomical units, AU) from the planet to the sun. How many Earth years would it take for a planet that is 6.76 AU from the sun? (1 point)15.23 17.58 154.46 3.58

OpenStudy (amistre64):

please close questions that you are done with so that you may post a new question in the "ask a question" box.

OpenStudy (anonymous):

OK